3. Part II Answer all questions. Show your solution in detail. Part credit is allowed. I Bing Co. produces a product that requires 3 gallons per unit. The Standard price is $18.50 per gallon. If 2500 units require 8000 gallons, which were purchased at $18.00 per gallon, calculate the direct materials (A) Price Variance, (B) Quantity Variance, and (C) Overall Cost Variance. (18 Pts.) Il Bing Co. produces a product that requires 2 standard hours of labor per unit at a standard hourly rate of $18 per hour. If 2500 units required 5500 hours of labor at a rate of $19 per hour, find the direct labor (A) Rate Variance, (B) Time Varlance, and () Overall Cost Variance. (18 Pts.)
Answer #1

Answer to Question 1:

Standard materials per unit = 3 gallons
Standard price per gallon = $18.50
Actual unit produced = 2,500
Actual materials used = 8,000 gallons
Actual price per gallon = $18.00

Standard materials allowed = Standard materials per unit * Actual unit produced
Standard materials allowed = 3 * 2,500
Standard materials allowed = 7,500 gallons

Direct Materials Price Variance = Actual materials used * (Actual price per gallon - Standard price per gallon)
Direct Materials Price Variance = 8,000 * ($18.00 - $18.50)
Direct Materials Price Variance = $4,000 Favorable

Direct Materials Quantity Variance = Standard price per gallon * (Actual materials used - Standard materials allowed)
Direct Materials Quantity Variance = $18.50 * (8,000 - 7,500)
Direct Materials Quantity Variance = $9,250 Unfavorable

Direct Materials Variance = Direct Materials Price Variance + Direct Materials Quantity Variance
Direct Materials Variance = $4,000 Favorable + $9,250 Unfavorable
Direct Materials Variance = $5,250 Unfavorable

Answer to Question 2:

Standard labor hours per unit = 2
Standard rate per hour = $18.00
Actual unit produced = 2,500
Actual labor hours = 5,500
Actual rate per hour = $19.00

Standard labor hours allowed = Standard labor hours per unit * Actual unit produced
Standard labor hours allowed = 2 * 2,500
Standard labor hours allowed = 5,000

Direct Labor Rate Variance = Actual labor hours * (Actual rate per hour - Standard rate per hour)
Direct Labor Rate Variance = 5,500 * ($19.00 - $18.00)
Direct Labor Rate Variance = $5,500 Unfavorable

Direct Labor Time Variance = Standard rate per hour * (Actual labor hours - Standard labor hours allowed)
Direct Labor Time Variance = $18.00 * (5,500 - 5,000)
Direct Labor Time Variance = $9,000 Unfavorable

Direct Labor Variance = Direct Labor Rate Variance + Direct Labor Time Variance
Direct Labor Variance = $5,500 Unfavorable + $9,000 Unfavorable
Direct Labor Variance = $14,500 Unfavorable